A box contains 14 large marbles and 10 small marbles. Each marble is either green or white. 8 of the large marbles are green, and 4 of the small marbles are white. If a marble is randomly selected from the box, what is the probability that it is large or white? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Number of marbles = 14+10 = 24

The information given in the question statement can be tabulated as below:

Large Small Total
Green 8 6 14
White 6 4 10
Total 14 10 24

Probability that the selected marble is large or white:

P(Large \cup White)=P(Large) +P(White)-P(Large \cap White)

\Rightarrow P(Large \cup White)=\frac{14}{24} +\frac{10}{24}-\frac{6}{24}=\frac{18}{24}=\frac{3}{4}=0.7500